# Towards the Worst-case Robustness of Large Language Models

### v3

@@ Abstract
Recent studies have revealed the vulnerability of Large Language Models (LLMs) to adversarial attacks, where adversaries craft specific inputs to induce wrong or even harmful outputs. Although various empirical defenses have been proposed, their worst-case robustness remains unexplored, raising concerns about the vulnerability to future stronger adversaries. In this paper, we systematically study the worst-case robustness of LLMs from both empirical and theoretical perspectives. First, we upper bound the worst-case robustness of deterministic defenses using enhanced white-box attacks, showing that most of them achieve nearly 0% robustness against white-box adversaries. Then, we derive a general tight lower bound for randomized smoothing using fractional or 0-1 knapsack solvers, and apply them to derive theoretical lower bounds of the worst-case robustness for previous stochastic defenses. For example, we certify the robustness of GPT-4o with uniform kernel smoothing against any possible attack, with an average $\ell_{0}$ perturbation of 2.02 or an average suffix length of 6.41 on the AdvBench dataset.
@@ 1 Introduction
Large Language Models (LLMs) [1, 2, 3] have gained significant attention in recent years due to their impressive performance in a wide range of applications, demonstrated substantial potential in both academic research and practical deployments, making them valuable assets in various domains [4, 5, 6, 7]. However, concerns about the adversarial robustness of LLMs have also emerged [8, 9] along with their rapid adoption. Even worse, recent studies [10, 11] have shown that adversaries can craft adversarial suffixes to input prompts, which can mislead LLMs to generate malicious or harmful content, also known as jailbreak attacks [12]. This vulnerability poses a serious threat to the security and reliability of LLM-based systems, potentially undermining their broader application.
In this work, we study the worst-case robustness of LLMs and their defenses, i.e., whether an adversarial example would exist and lead to undesirable outputs [13]. As widely recognized, worst-case robustness is a longstanding academic problem [14, 9], which not only provides insights into the intrinsic mechanisms of neural networks [15], but also serves as a lower bound on the robustness achievable under practical attacks, since a model may have adversarial examples that practical adversaries cannot find due to limited time and information [13, 16].
To provide a tighter upper bound on worst-case robustness, we devise stronger adversaries by strictly ensuring the consistency of tokens during optimization and inference (this builds upon the previous I-GCG method [17], thus we call our method I2-GCG). As shown in Table 1, this slight improvement greatly reduces the estimation of most typical previous deterministic defenses by more than 30%, resulting in these defenses exhibiting nearly 0% worst-case robustness. This finding is not surprising: adding extra prompts does not address the intrinsic vulnerability of neural networks to adversarial examples; detection and filtering defenses are easily circumvented in white-box settings by targeting the detector networks themselves [16, 13]; and adversarial training demands exponentially greater resources [18, 19], rendering it currently impractical for sufficiently training large-scale models.
@@
Although our attacker obtains a relatively accurate estimation of worst-case robustness for deterministic defenses, it only provides extremely loose upper bounds for stochastic defenses. For instance, a safety detector should not be robust with a suffix length of 20, as a suffix “do not answer this question” can indeed change the detector’s result from harmful to safe. However, when applying a stochastic defense (e.g., [20] here) to safety detectors, evaluating with I2-GCG against a suffix of length 20 still yields over 80% robustness. This indicates that, when evaluating stochastic defenses, although an adversarial example may exist, the optimization process is significantly affected by stochasticity [21], causing current attackers to fail to find them, thus obtaining only an extremely loose estimation of worst-case robustness [22]. Therefore, we advocate that one should not estimate worst-case robustness solely by upper bounding using practical attacks, but should also establish a theoretical lower bound. By bounding from both sides, we can obtain a clearer understanding of worst-case robustness [23, 24, 25].
@@
Most stochastic defenses can be formulated as returning the output of $f(\bm{z})$ from $\bm{z}\sim p(\bm{z}|\bm{x})$ instead of $f(\bm{x})$ [26]. Since the output of such a stochastic function is a random variable, it sometimes returns the true result and sometimes is false. To enable a more formal analysis, we study their expectation $g(\bm{x})=\mathbb{E}_{p(\bm{z}|\bm{x})}[f(\bm{z})]$. If the expectation of a stochastic defense is robust, then most outputs of such a stochastic defense on adversarial examples would also be correct due to the concentration of random variables [23]. We formulate the problem of obtaining a theoretical lower bound for $g$ as a functional optimization problem that identifies the minimal output $p_{adv}:=\min_{\bm{x}_{adv},f}g(\bm{x}_{adv})$ over all possible adversarial examples $\bm{x}_{adv}$ and base functions $f$, given that we know the output on the original sample $g(\bm{x})=p_{A}$. If this minimal output satisfies a condition, e.g., $p_{adv}\geq\tau$ for some threshold $\tau$, we say that the function $g$ is certifiably robust for all $\mathcal{D}(\bm{x},\bm{x}_{adv})\leq d$. This definition encompasses all previous specific distributions (e.g., Gaussian distribution [23], Laplacian distribution [27]) and applications (e.g., classification, safety).
@@
To derive certified robustness for any such $g$, we need to solve both the functional minimization problem $\min_{f}$ and the input minimization problem $\min_{\bm{x}_{adv}}$. We first show that the functional minimization problem $\min_{f}$ can be reduced to the Fractional Knapsack problem when $f$ is a bounded function, or the 0-1 Knapsack problem when $f$ is a binary function, with the capacity of knapsack $p_{A}:=g(\bm{x})$, the value of each item $-p(\bm{z}|\bm{x}_{adv})$, and the weight of each item $p(\bm{z}|\bm{x})$. This only slightly differs from the standard knapsack problem, which requires the total weight of items to be less than or equal to the capacity (i.e., $g(\bm{x})\leq p_{A}$), whereas here we require $g(\bm{x})=p_{A}$. This can be addressed by slightly modifying the greedy algorithm for the Fractional Knapsack problem and the dynamic programming approach for the 0-1 Knapsack problem. To solve the input minimization problem $\min_{\bm{x}_{adv}}$, one can simply enumerate all $\bm{x}_{adv}$ and run the knapsack solver for each $\bm{x}_{adv}$, identifying the minimal one as the lower bound. In the main text, we show that it is not necessary to enumerate all $\bm{x}_{adv}$ due to an equivalence: if this $f$ performs worst on a given $\bm{x}_{adv}$, then there exists another $f$ that performs worst on a different $\bm{x}_{adv}$. In other words, the functional minimization $\min_{f}$ already implicitly solves most of the input minimization $\min_{\bm{x}_{adv}}$.
@@
Based on these solvers, we provide theoretical lower bounds for several previous empirical defenses, including random masking [28, 29], random perturbation on tokens [20, 30], random perturbation of characters [31]. We present these results on Table 2 and Table 3. For example, we certify the robustness of a specific case, i.e., smoothing the Llama-3 safety detector [3] using a uniform kernel [30, 20], against any possible attack with an average $\ell_{0}$ perturbation of 1.54 or an average suffix length of 6.41 on the AdvBench dataset.
@@ 2 Backgrounds and Preliminaries
Worst-case robustness, white-box attacks, and practical attacks. The existence of adversarial examples [15] is a long-standing challenge in the safety of deep learning models. By adding small, often imperceptible perturbations to the input, deep learning models can be misled into producing incorrect outputs [32]. Worst-case robustness is defined as whether there exist adversarial examples within a specified neighborhood of normal examples [15, 33]. Thus, it serves as a lower bound on the robustness achievable under attacks, since a model may have adversarial examples that optimizers cannot find [16]. White-box robustness is defined as robustness against white-box adaptive attacks, where the attacker has full access to the model and defense strategies, thereby providing an upper bound estimation for worst-case robustness [13]. Black-box robustness refers to robustness against attackers with certain constraints, e.g., limited access to the gradient [13], limited time [34]. Clearly, black-box robustness also serves as an upper bound for white-box robustness [16]. Evaluating worst-case robustness provides a lower bound on robustness against potential real-world threats [35] and helps us understand the intrinsic mechanisms of neural networks [15, 33].
@@
Jailbreaking attacks and defenses. Recently, jailbreaking attacks have emerged as a specific type of adversarial attack designed to manipulate LLMs into generating harmful, violent, or private content misaligned with human values. These attacks pose a significant safety concern for the deployment of LLMs [10]. One category of jailbreaking attacks employs heuristic methods, such as manually crafted prompts [36, 37], or utilizes LLMs to generate jailbreaking prompts [11, 38]. Another category uses optimization-based methods, which minimize a formulated jailbreaking loss to generate adversarial prompts [10, 17, 39]. In this work, we focus on the latter approach, as it can be mathematically formulated and analyzed. To address the safety concerns posed by jailbreaking, various defenses have been proposed, including prompt detection [40], adversarial training [41], and additional safety prompts [42]. However, these defenses primarily target black-box attacks. When evaluated under stronger white-box attacks, most of the deterministic defenses exhibit nearly 0% robustness (detailed in Section 3).
@@
Certified robustness. Neural networks are generally composed of multiple stacked linear layers. Their maximum Lipschitz is approximately the product of the maximum singular values of these linear layers, which can be sufficiently large [43]. As a result, even small perturbations in the input can significantly alter their outputs [33]. Directly verifying whether a function is robust for all $\bm{x}_{adv}$ would require enumerating all possible inputs, which is NP-hard [44]. To address this challenge, researchers propose randomized smoothing [23, 45], which constructs a smoothed function $g$ by aggregating the ensemble predictions of a base function $f$ over a perturbation distribution $p(\bm{z}|\bm{x})$ by $g(\bm{x})=\mathbb{E}_{p(\bm{z}|\bm{x})}[f(\bm{z})]$. Thanks to the mathematical properties of the smoothed function $g$, it exhibits inherent smoothness regardless of the vulnerability of the base function $f$. For instance, Cohen et al. [23] demonstrate that when $p(\bm{z}|\bm{x})=\mathcal{N}(\bm{0},\bm{I})$, the resulting smoothed function $g$ is guaranteed to be at least $\frac{1}{\sqrt{2\pi}}$-Lipschitz, independent of how susceptible $f$ is to adversarial perturbations. Therefore, if we know $g(\bm{x})=p_{A}$, then we can show that $g(\bm{x}_{adv})\geq p_{A}-\frac{1}{\sqrt{2\pi}}$ for all $\|\bm{x}_{adv}-\bm{x}\|_{2}\leq 1$.
@@ 3 Upper Bounding Worst-case Robustness
Motivation. As suggested by Carlini et al. [9], there are two primary reasons researchers focus on worst-case robustness. On the one hand, worst-case robustness represents the maximum capability of real adversaries. If our model achieves reasonable worst-case robustness, we can guarantee that it is safe against any adversaries [13]. On the other hand, worst-case robustness provides insight into the worst-case behavior of a neural network, even if we do not believe real adversaries can achieve such worst-case [46]. Understanding worst-case robustness helps us gain a deeper understanding of the intrinsic mechanisms of neural networks [15].
White-box robustness, where the attacker has full knowledge of the defended model, represents an upper bound for the worst-case robustness. The actual worst-case robustness must be smaller than the robustness achieved by a white-box attacker [47]. Conversely, white-box robustness serves as a lower bound of robustness that an attacker can achieve in practical scenarios, such as black-box settings, where the attacker has limited access to the model’s internal parameters. Therefore, it helps identify vulnerabilities that might be exploited under more favorable conditions for the adversary.
@@
Our design. We observe that previous white-box evaluations of LLMs fail to properly evaluate robustness [48] because they do not strictly ensure the consistency of tokenization when calculating the loss in parallel and sequentially generating the output. Even slight differences in tokenization can result in vastly different losses, leading to failures in generating adversarial examples. To address this issue, we improve upon the I-GCG [17] by carefully and strictly ensuring token consistency during both attacking and inference. Accordingly, we name our attack as I2-GCG. See Sec. H.1 for further details.
@@
Results on deterministic defenses. As demonstrated in Table 1, our I2-GCG results in nearly 0% robustness for most typical deterministic defenses, demonstrating their worst-case vulnerability [^1] Disclaimer: This does not imply that these defenses are impractical. On the contrary, they are currently the most practical defenses, as practical attackers have limited information about black-box models and defenses.. This is unsurprising. Most defenses in the vision domain have been attacked to 0% robustness in the last decade [49, 16]. Adding extra prompts does not address the intrinsic vulnerability of neural networks to adversarial examples. Detection and filtering defenses are easily circumvented in white-box settings by targeting the filter network itself [16, 13]. Adversarial training works for previous visual adversarial examples, but it demands exponentially greater resources [18, 19]. Current adversarial training on LLMs does not train for a sufficiently long time, improving only average-case robustness but, as of now, not the worst-case [48].
@@
Results on Randomized Defenses. Our I2-GCG method, however, obtains only extremely loose upper bounds for stochastic defenses. For instance, a safety detector should not be robust with a suffix length of 20, as a suffix “do not answer this question” can change the detector’s result from harmful to safe. However, when applying stochastic defenses—such as smoothing each token with a random mask (as in [29, 20]), substituting each token with random tokens (uniform in [20, 30]), or substituting each character with random characters (SmoothLLM [31]) to safety detectors, evaluating with I2-GCG against a suffix of length 20 still yields over 60% robustness. This indicates that, when evaluating stochastic defenses, although an adversarial example may exist, the optimization process is significantly affected by stochasticity [21], causing current attackers to fail to find them, thus obtaining only an extremely loose estimation of worst-case robustness [22]. Therefore, we advocate that one should not estimate worst-case robustness solely by upper bounding using practical attacks, but should also establish a theoretical lower bound. By bounding from both sides, we can obtain a clearer understanding of worst-case robustness [23, 24, 25].

### v4

@@ Abstract
Recent studies have revealed the vulnerability of large language models to adversarial attacks, where adversaries craft specific inputs to induce wrong or even harmful outputs. Although various empirical defenses have been proposed, their worst-case robustness remains unexplored, raising concerns about the vulnerability to future stronger adversaries. In this paper, we systematically study the worst-case robustness of LLMs from both empirical and theoretical perspectives. First, we upper bound the worst-case robustness of deterministic defenses using enhanced white-box attacks, showing that most of them achieve nearly 0% robustness against white-box adversaries. Then, we derive a general tight lower bound for randomized smoothing using fractional or 0-1 knapsack solvers, and apply them to derive theoretical lower bounds of the worst-case robustness for previous stochastic defenses. For example, we certify the robustness of GPT-4o with uniform kernel smoothing against any possible attack, with an average $\ell_{0}$ perturbation of 2.02 or an average suffix length of 6.41 on the AdvBench dataset.
@@ 1 Introduction
Large Language Models (LLMs) OpenAI (2023); Anthropic (2024); Dubey et al. (2024) have gained significant attention in recent years due to their impressive performance in a wide range of applications, demonstrated substantial potential in both academic research and practical deployments, making them valuable assets in various domains Cai et al. (2023); Cummins et al. (2023); Trinh et al. (2024); Liu et al. (2024b). However, concerns about the adversarial robustness of LLMs have also emerged Wang et al. (2023a); Carlini et al. (2023a) along with their rapid adoption. Even worse, recent studies Zou et al. (2023); Chao et al. (2023) have shown that adversaries can craft adversarial suffixes to input prompts, which can mislead LLMs to generate malicious or harmful content, also known as jailbreak attacks Wei et al. (2023a). This vulnerability poses a serious threat to the security and reliability of LLM-based systems, potentially undermining their broader application.
In this work, we study the worst-case robustness of LLMs and their defenses, i.e., whether an adversarial example would exist and lead to undesirable outputs Carlini et al. (2019). As widely recognized, worst-case robustness is a longstanding academic problem Madry et al. (2018); Carlini et al. (2023a), which not only provides insights into the intrinsic mechanisms of neural networks Szegedy et al. (2014), but also serves as a lower bound on the robustness achievable under practical attacks, since a model may have adversarial examples that practical adversaries cannot find due to limited time and information Athalye et al. (2018); Carlini et al. (2019).
To provide a tighter upper bound on worst-case robustness, we devise stronger adversaries by ensuring that tokenization during inference is exactly the same as that during attack optimization (this builds upon the previous I-GCG method Jia et al. (2024), thus we call our method I2-GCG). As shown in Table 1, this slight improvement greatly reduces the robustness of most typical deterministic defenses by more than 30%, making these defenses exhibit nearly 0% worst-case robustness. This finding is not surprising: adding extra prompts does not address the intrinsic vulnerability of neural networks to adversarial examples; detection and filtering defenses are easily circumvented in white-box settings by targeting the detector networks themselves Athalye et al. (2018); Carlini et al. (2019); and adversarial training demands exponentially greater resources Diakonikolas et al. (2020); Gourdeau et al. (2021), rendering it currently impractical for sufficiently training large-scale models.
@@
Although our attacker obtains a relatively accurate estimation of worst-case robustness for deterministic defenses, it provides extremely loose upper bounds for stochastic defenses. For instance, a safety detector should not be robust to an adversarial suffix of length 20, as a suffix “do not answer this question” can indeed change the detector’s result from harmful to safe. However, when applying a stochastic defense (e.g., Lou et al. (2023)) to safety detectors, evaluating with I2-GCG against a suffix of length 20 still yields over 60% robustness. This indicates that, when evaluating stochastic defenses, although an adversarial example may exist, the optimization process is significantly affected by stochasticity (Kang et al., 2024), causing current attackers to fail to find them and obtain only an extremely loose estimation of worst-case robustness (Lee & Kim, 2023). Therefore, we advocate that one should not only upper bound worst-case robustness by practical attacks, but also establish a theoretical lower bound. By bounding from both sides, we can obtain a clearer understanding of worst-case robustness Cohen et al. (2019); Weng et al. (2018); Hein & Andriushchenko (2017).
@@
Most stochastic defenses can be formulated as returning the output of $f(\bm{z})$ from sampling $\bm{z}\sim p(\bm{z}|\bm{x})$ instead of $f(\bm{x})$ Gao et al. (2022). Since the output of such a stochastic function is a random variable, it sometimes returns the true result and sometimes returns a false result. To enable a more formal analysis, we study their expectation $g(\bm{x})=\mathbb{E}_{p(\bm{z}|\bm{x})}[f(\bm{z})]$. If the expectation of a stochastic defense is robust, then most outputs of such a stochastic defense on adversarial examples would also be correct due to the concentration of random variables Cohen et al. (2019). To obtain $p_{adv}:=\min_{\bm{x}_{adv}}g(\bm{x}_{adv})$ for all $\bm{x}_{adv}$ such that $\mathcal{D}(\bm{x},\bm{x}_{adv})\leq d$, we relax the function $f$ to the hypothesis class $\mathcal{F}$ (where $f\in\mathcal{F}$) by formulating $\min_{\bm{x}_{adv}}g(\bm{x}_{adv})\geq\min_{\bm{x}_{adv}}\min_{f^{\prime}\in\mathcal{F}}\sum_{\bm{z}}f^{\prime}(\bm{z})p(\bm{z}|\bm{x}_{adv})$. This relaxation introduces symmetrization, such that solving $\min_{f^{\prime}\in\mathcal{F}}$ typically yields the result for $\min_{\bm{x}_{adv}}$, as the worst-case function’s output of these inputs are equivalent (see Sec.˜4.2 for details).
@@
Therefore, to obtain the lower bound for $\min_{\bm{x}_{adv}}g(\bm{x}_{adv})$, we only need to solve the functional minimization problem $\min_{f^{\prime}}$ instead of the input minimization problem $\min_{\bm{x}_{adv}}$. We show that the functional minimization problem $\min_{f^{\prime}}$ can be reduced to the Fractional Knapsack problem when $f$ is a bounded function, or to the 0-1 Knapsack problem when $f$ is a binary function, with the knapsack capacity $p_{A}:=g(\bm{x})$, the value of each item as $p(\bm{z}|\bm{x}_{adv})$, and the weight of each item as $p(\bm{z}|\bm{x})$. This differs slightly from the standard knapsack problem, which requires the total weight of items to be less than or equal to the capacity (i.e., $g(\bm{x})\leq p_{A}$), whereas we require $g(\bm{x})=p_{A}$. This constraint can be addressed by slightly modifying the greedy algorithm for the Fractional Knapsack problem and the dynamic programming approach for the 0-1 Knapsack problem. Note that our bound is black-box tight, i.e., if $g(\bm{x})=p_{A}$ is the only known information, it is impossible to obtain a higher $\min_{\bm{x}_{adv}}g(\bm{x}_{adv})$ than that provided by knapsack solvers. The results of fractional knapsack solvers are also equivalent to prior results in specific distributions, e.g., Gaussian distributions Cohen et al. (2019), Laplace distributions Teng et al. (2020).
@@
Based on these solvers, we provide theoretical lower bounds for several previous empirical defenses, including random masking Ye et al. (2020); Zeng et al. (2023), random perturbation on tokens Lou et al. (2023), and on characters Robey et al. (2023). We present the results in Table˜3 and Table˜3. For example, we certify the robustness of a specific case, i.e., smoothing the GPT-4o safety detector using a uniform kernel Lou et al. (2023), against any possible attack, with an average $\ell_{0}$ perturbation of 2.02 or an average suffix length of 6.41 on the AdvBench dataset.
@@ 2 Backgrounds and Preliminaries
Worst-case robustness, white-box attacks, and practical attacks. Adversarial examples Szegedy et al. (2014) is a long-standing problem for the safety of deep learning models. Worst-case robustness is defined as whether there exist adversarial examples within a specified neighborhood of normal examples Carlini & Wagner (2017b). Thus, it serves as a lower bound on the robustness achievable under attacks, since a model may have adversarial examples that optimizers cannot find Athalye et al. (2018). White-box robustness is defined as robustness against white-box adaptive attacks, where the attacker has full access to the model and defense strategies, thereby providing an upper bound estimation for worst-case robustness Carlini et al. (2019). Black-box robustness refers to robustness against attackers with certain constraints, e.g., limited access to the gradient Carlini et al. (2019), limited time Papernot et al. (2017). Evaluating worst-case robustness provides a lower bound against potential real-world threats Croce & Hein (2020) and helps us understand the intrinsic mechanisms of neural networks Szegedy et al. (2014); Goodfellow et al. (2015).
@@
Jailbreaking attacks and defenses. Recently, jailbreaking attacks have emerged as a specific type of adversarial attack to manipulate LLMs into generating harmful, violent, or private content misaligned with human values. These attacks pose a significant safety concern for the deployment of LLMs Zou et al. (2023). One category of jailbreaking attacks employs heuristic methods, such as manually crafted prompts Wei et al. (2023b); Jailbreak Chat (2024), or utilizes LLMs to generate jailbreaking prompts Chao et al. (2023); Mehrotra et al. (2023). Another category uses optimization-based methods, which minimize a formulated jailbreaking loss to generate adversarial prompts Zou et al. (2023); Jia et al. (2024); Liu et al. (2023). In this work, we focus on the latter approach, as it can be mathematically formulated and analyzed. To address the safety concerns posed by jailbreaking, various defenses have been proposed, including prompt detection Alon & Kamfonas (2023), adversarial training Mo et al. (2024), and additional safety prompts Wu et al. (2023). However, these defenses primarily target black-box attacks. When evaluated under stronger white-box attacks, most of the deterministic defenses exhibit nearly 0% robustness (detailed in Section 3).
@@
Certified robustness. Neural networks are generally composed of multiple stacked linear layers. Their maximum Lipschitz is approximately the product of the maximum singular values of these linear layers, which can be sufficiently large Fazlyab et al. (2019). As a result, even small perturbations in the input can significantly alter their outputs Goodfellow et al. (2015). Verifying ReLU networks has been shown to be NP-complete Katz et al. (2017), and they lack efficient approximation algorithms in the worst case Weng et al. (2018), making them challenging to scale to large models. To address this challenge, researchers propose randomized smoothing Cohen et al. (2019); Salman et al. (2019), which constructs a smoothed function $g$ by aggregating the ensemble predictions of a base function $f$ over a perturbation distribution $p(\bm{z}|\bm{x})$ by $g(\bm{x})=\mathbb{E}_{p(\bm{z}|\bm{x})}[f(\bm{z})]$. Thanks to the mathematical properties of the smoothed function $g$, it exhibits inherent smoothness regardless of the vulnerability of the base function $f$. For instance, Cohen et al. (2019) demonstrate that when $p(\bm{z}|\bm{x})=\mathcal{N}(\bm{0},\bm{I})$, the resulting smoothed function $g$ is guaranteed to be at least $\frac{1}{\sqrt{2\pi}}$-Lipschitz, independent of how susceptible $f$ is to adversarial perturbations. Therefore, if we know $g(\bm{x})=p_{A}$, then we can show that $g(\bm{x}_{adv})\geq p_{A}-\frac{1}{\sqrt{2\pi}}$ for all $\|\bm{x}_{adv}-\bm{x}\|_{2}\leq 1$.
@@ 3 Upper Bounding Worst-case Robustness
Following common practice Carlini & Wagner (2017a), we use white-box attacks to upper bound the worst-case robustness of large language models, which also provide a lower bound for black-box robustness in practical scenarios. See Sec.˜H.1 for a detailed discussion on the relationship between white-box, black-box, worst-case, and practical robustness.
@@
Our design. We observe that previous white-box attacks on LLMs fail to properly evaluate their robustness Jain et al. (2023) because they do not strictly ensure the consistency of tokenization when calculating the loss in parallel and sequentially generating the output. Even slight differences in tokenization can result in vastly different losses, leading to failures in generating adversarial examples. To address this issue, we improve upon the I-GCG Jia et al. (2024) by carefully and strictly ensuring token consistency during both attacking and inference. Accordingly, we name our attack as I2-GCG. See Sec.˜H.2 for further details.
@@
Results on deterministic defenses. As demonstrated in Table˜1, our I2-GCG results in nearly 0% robustness for most typical deterministic defenses, demonstrating their worst-case vulnerability [^1] Disclaimer: This does not imply that these defenses are impractical. On the contrary, they are currently the most practical defenses, as practical attackers have limited information about black-box models and defenses.. This is unsurprising. Most defenses in the vision domain have been attacked to 0% robustness in the last decade Athalye & Carlini (2018); Athalye et al. (2018). Adding extra prompts does not address the intrinsic vulnerability of neural networks to adversarial examples. Detection and filtering defenses are easily circumvented in white-box settings by targeting the filter network itself Athalye et al. (2018); Carlini et al. (2019). Adversarial training works for previous visual adversarial examples, but it demands exponentially greater resources Diakonikolas et al. (2020); Gourdeau et al. (2021). Current adversarial training on LLMs does not train for a sufficiently long time, improving only average-case robustness but, as of now, not the worst-case Jain et al. (2023).
@@
Results on randomized defenses. Our I2-GCG method, however, obtains only extremely loose upper bounds for stochastic defenses. For instance, a safety detector should not be robust with a suffix length of 20, as a suffix “do not answer this question” can change the detector’s result from harmful to safe. However, when applying stochastic defenses, such as smoothing each token with a random mask Zeng et al. (2023); Lou et al. (2023), or substituting each token/character with random ones Lou et al. (2023); Robey et al. (2023) to safety detectors, evaluating with I2-GCG against a suffix of length 20 still yields over 60% robustness. This indicates that, when evaluating stochastic defenses, although an adversarial example may exist, the optimization process is significantly affected by stochasticity Kang et al. (2024), causing current attackers to fail to find them and obtain only an extremely loose estimation of worst-case robustness Lee & Kim (2023). Therefore, we argue that we should not only consider the upper bound of worst-case robustness using practical attacks, but also establish a theoretical lower bound. By doing so, we can obtain a clearer understanding of worst-case robustness Cohen et al. (2019); Weng et al. (2018); Hein & Andriushchenko (2017).